OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

J = 10 + t t - 3 = 2J

OpenStudy (anonymous):

Solve for t

OpenStudy (anonymous):

with substitution

OpenStudy (anonymous):

is it 13?

OpenStudy (anonymous):

J = 10 + t t = 2J + 3 substitute t in for J you get t = 2(10+t) + 3

OpenStudy (anonymous):

hmm.. wait that's wrong.

OpenStudy (anonymous):

well if you put in 13

OpenStudy (anonymous):

t+10=j j-3=2(t-3)

OpenStudy (anonymous):

Yes ^ That's what i meant =P

OpenStudy (anonymous):

so whts the answer would be?

OpenStudy (anonymous):

Lol completeidiot saving me all of the times when my brain malfunctions. Thx

OpenStudy (anonymous):

substituting for j t+10-3=2(t-3) solve for t yourself

OpenStudy (anonymous):

Yes. Your answer is right.

OpenStudy (anonymous):

my equation is correct, i have not given an answer

OpenStudy (anonymous):

13?

OpenStudy (anonymous):

yea look about right

OpenStudy (anonymous):

now back to loitering around chat and being bored